OpenStudy (anonymous):

explain what happens to y=f(x) if y=-1/2f(x)

OpenStudy (shubhamsrg):

any attempts/guesses ?

OpenStudy (anonymous):

y=-\[\frac{ 1 }{ 2 }\]f(x)

OpenStudy (anonymous):

I think it's a verticle expansion by a factor of 2

OpenStudy (shubhamsrg):

thats one of the 4 choices you have right ?

OpenStudy (anonymous):

it's' not multiple choice. I'm supposed to explain what happens to y=f(x)

OpenStudy (shubhamsrg):

hmm. Well, y=-1/2 f(x) would mean the height gets reflected on the x axis , and also are reduced to 1/2 .

OpenStudy (shubhamsrg):

I mean 1/2 of earlier height.

OpenStudy (anonymous):

so a verticle compression by a factor of half, and it flips upside down? or rightside left?

OpenStudy (shubhamsrg):

you should be able to relate it from the fact that y co-ordinate becomes negative if it was positive and vice-versa if y co-ordinate was negative.

OpenStudy (anonymous):

can you just give me the answer. it's an open-book quiz, and i have to go! thank you so much

OpenStudy (shubhamsrg):

no I can not just give you the answer. and you are welcome.

OpenStudy (anonymous):

sure you can! it's open-book. That means that we can get the answer by whatever means necessary. It's not a test...

OpenStudy (anonymous):

I will give you 'best response'

OpenStudy (shubhamsrg):

lol.. :D

OpenStudy (anonymous):

plus, i think i figured it out, just tell me if i'm right

OpenStudy (anonymous):

it's a verticle compression by a factor of 1/2, and a reflection in the y axis